OpenStudy (anonymous):

Write the equation of the vertical asymptote for the function, y = ln (x - 1)

ganeshie8 (ganeshie8):

y = ln(x-1) is a graph u get by shifting its parent function y = ln(x) to right by 1 unit

ganeshie8 (ganeshie8):

do u knw whats the vertical asymptote of ln(x) ?

OpenStudy (anonymous):

No, I don't

ganeshie8 (ganeshie8):

okay, vertical asymptote of ln(x) is x=0

ganeshie8 (ganeshie8):

since ln(x-1) is shifted version of ln(x) to right by 1unit, the corresponding vertical asymptote also shifts to right by 1 unit.

ganeshie8 (ganeshie8):

so the vertical asymptote of ln(x-1) wud be x = 1

OpenStudy (anonymous):

okay, thank you so much for that :)

ganeshie8 (ganeshie8):

np :)